### v2

# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

Abstract
A major bottleneck in uplink distributed massive multiple-input multiple-output networks is the sub-optimal performance of local combining schemes, coupled with high fronthaul load and computational cost inherent in centralized large scale fading decoding (LSFD) architectures. This paper introduces a decentralized decoding architecture that fundamentally breaks from the LSFD, by allowing each access point (AP) to calculate interference-suppressing local weights independently and apply them to its data estimates before transmission. Furthermore, two generalized local zero-forcing (ZF) frameworks, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each AP adaptively and independently determines its combining strategy through a local sum spectral efficiency (SE) optimization that classifies user equipments (UEs) as strong or weak, eliminating the fixed thresholds used in the PFZF and PWPFZF schemes. To enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot. The closed-form SE expressions corresponding to the proposed schemes are derived. Numerical results show that the proposed schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overheads and computation costs with lower performance penalty compared to them.
I Introduction
Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee uniform service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2]. To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [3, 4]. A major challenge in this regard is the development of effective uplink combining schemes that can alleviate inter-user interference within practical limitations.
Centralized zero-forcing (ZF) combining, as studied in [5, 6, 7], requires substantial fronthaul signaling and imposes prohibitively high computational costs. Similarly, centralized minimum mean square error (MMSE) combining [8] suffers from both high computation cost and fronthaul overhead, which makes it impractical for large-scale deployments. To alleviate this, partial MMSE combining is proposed in [9, 3, 10], where the interference from only a subset of UEs is considered when designing the combining vector for each UE. Along similar lines, partial centralized approaches have been explored in [11, 12], where joint ZF is applied across a subset of APs based on stronger channels. These partial centralized ZF schemes reduce the fronthaul load compared to fully centralized ZF, but the overhead still scales with the number of UEs, as for each UE, a subset of APs need to send the channel estimates to the CPU, which limits their suitability for ultra-dense network scenarios. These centralized and partial centralized combining schemes follow the fully centralized architecture as shown in Fig. 1, which has very high computational cost and fronthaul overhead, as for all UEs, the channel needs to be estimated at the CPU with the cooperation of all or a subset of APs.
Several distributed combining approaches have been proposed in the literature. Maximum ratio (MR) combining, introduced in [1], is computationally simple and fronthaul-efficient, but its performance is fundamentally limited by its inability to effectively mitigate inter-user interference. To improve upon the MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance. However, a key drawback of this approach is that closed-form expressions for the spectral efficiency (SE) cannot be derived. As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult. In [1, 3], the authors follow a simple decoding architecture as in Fig. 1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality with respect to centralized decoding.
The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form SE expressions. In these schemes, each AP classifies the UEs into strong and weak groups, generates one combining vector per orthogonal pilot: the ZF for strong UEs and the MR for weak UEs, and assigns scaled versions to individual UEs based on their estimated channel variances [13]. The fundamental limitations of the PFZF and PWPFZF schemes stem from two factors. First, the rigid, threshold-based user grouping may misclassify UEs, leading to suboptimal interference suppression and inefficient use of spatial resources. Second, all APs are constrained to follow the same combining scheme, which prevents each AP from independently adapting its strategy to local channel conditions and interference patterns, further limiting overall system performance.
The large-scale fading decoding (LSFD) as shown in Fig. 1, is employed at the CPU to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14]. In the optimal LSFD approach [13], each AP must transmit both channel and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16]. To alleviate this burden, partial LSFD is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process. However, partial LSFD still requires matrix inversions of the same dimension as optimal LSFD and transmission of the channel estimates to the CPU. As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of the LSFD-based architectures in ultra-dense D-mMIMO deployments.
(a) FullyCentralized(b) LSFD(c) SimpleDecoding(d) Proposed DecentralizedDecoding APs CPU All Data Global Combining Global Estimates APs Local Combining CPU Local estimates $\hat{x}_{mt}$Channel estimates LSFD Weights Final Estimates APs Local Combining CPU $\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}\hat{x}_{mt}$ APs Local Combining Local Weights $a_{mt}$ CPU $a_{mt}\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}a_{mt}\hat{x}_{mt}$
Figure 1: Uplink processing architectures: (a) Fully centralized, (b) LSFD with coordination, (c) Simple decoding, (d) Proposed decentralized decoding.
I-A Contributions
This work addresses some of the above mentioned issues. Our principal contributions are as follows:
Decentralized Decoding Architecture with Local Weighting: We propose a novel uplink processing architecture as shown in Fig. 1, that completely eliminates the need for centralized LSFD and its associated fronthaul overhead. Central to this architecture is the design of interference-suppressing local weights, computed independently by each AP using only its local channel information. These weights are applied to data estimates before transmission, enabling the central processor to operate as a simple aggregator rather than a coordinator. This eliminates the need for sharing any channel information with the CPU. This transition from a coordinated to a truly decentralized architecture is the key enabler for the practical, large-scale deployment.
Adaptive Combining Strategy: We introduce generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF). Unlike the conventional PFZF and PWPFZF schemes, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy. Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR, thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
Novel Decentralized Local Optimization for Pilot Partitioning: We also propose novel decentralized optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum SE metric. This pilot-level optimization is solved using projected gradient ascent optimization, maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable SE of the proposed system. This analysis is essential due to the new statistical properties introduced by pilot-dependent processing and local weighting, and it provides critical insights for system design without the need for Monte Carlo simulation as in MMSE based schemes.
Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good SE with reduced computation cost with the state-of-art baselines. The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul overhead and computational costs in comparison to the LSFD framework and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
Organization: The remainder of the paper is organized as follows. Section II describes the system model, along with the channel estimation and data detection procedures. Section III presents the proposed combining schemes and derives their corresponding closed-form SE expressions. Section IV provides a rigorous performance evaluation. Finally, Section V concludes the paper and outlines potential directions for future work.
Notation: Scalars are denoted in italics (e.g., $x$), vectors by bold lowercase letters (e.g., $\mathbf{x}$), and matrices by bold uppercase letters (e.g., $\mathbf{X}$). The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate of a scalar is written as $(\cdot)^{*}$. The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$. The expectation operator is denoted by $\mathbb{E}[\cdot]$. The identity matrix of size $N$ is denoted by $\mathbf{I}_{N}$. A complex Gaussian random vector $\mathbf{x}$ with mean $\bm{\mu}$ and covariance matrix $\mathbf{K}$ is denoted as $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\mathbf{K})$. Also, $_{l}$ represent the $l$-th element of the vector $[.]$.
II System Model
The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas. Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs work together to service the UEs using the same frequency and time resources. Every AP is linked to the CPU by a fronthaul link, so that the network can operate together and process signals. The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for pilot training. The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$. The channel vector between AP $m$ and UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$. We suppose that $\textbf{h}_{mt}\sim\mathcal{CN}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{CN}(0,\beta_{mt}\textbf{I}_{A})$.
II-A Channel Estimation
A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of UE $t$. At the $m$-th AP, the received signal is:
$\textbf{y}^{pilot}_{m}=\sum\nolimits_{t=1}^{T}\sqrt{p^{p}_{t}L_{p}}\textbf{g}_{mt}\bm{\psi}_{i_{t}}^{{H}}+\textbf{N}_{m},$
where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by independent and identically distributed complex Gaussian entries. Additionally, $p^{p}_{t}$ represents the normalized pilot power for UE $t$ during pilot transmission. The MMSE estimate of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
$\hat{\textbf{g}}_{mt}=\sqrt{p^{p}_{t}L_{p}}\beta_{mt}\theta^{-1}_{mi_{t}}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
where $\theta_{mi_{t}}{=}\big(\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1\big)$. The estimate $\hat{\mathbf{g}}_{mt}$ and estimated error $\tilde{\mathbf{g}}_{mt}{=}{\mathbf{g}}_{mt}{-}\hat{\mathbf{g}}_{mt}$ are independent Gaussian with distributions $\hat{\mathbf{g}}_{mt}{\sim}\mathcal{CN}(0,\gamma_{mt}\textbf{I}_{A})$ and $\tilde{\mathbf{g}}_{mt}{\sim}\mathcal{CN}(0,(\beta_{mt}{-}\gamma_{mt})\textbf{I}_{A})$, where
$\gamma_{mt}=\mathbb{E}\big\{|[\hat{\mathbf{g}}_{mt}]_{l}|^{2}\big\}=p^{p}_{t}L_{p}\beta_{mt}^{2}\theta^{-1}_{mi_{t}},$ (1)
Remark 1: The channel estimates of UEs sharing the same pilot sequence are linearly dependent. As a result, an AP cannot fully distinguish among these UEs, leading to residual pilot contamination.
The matrix of estimated channels at AP $m$ is denoted by $\mathbf{\hat{G}}_{m}{=}[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]{\in}\mathbb{C}^{A\times T}$. Owing to pilot reuse among the UEs, the estimated channel vectors $\hat{\mathbf{g}}_{mt}$ and $\hat{\mathbf{g}}_{mk}$ corresponding to two UEs $t$ and $k$ that share the same pilot sequence become linearly dependent. As a result, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient. The full rank matrix is designed by removing the linearly dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence. Thus, the matrix $\mathbf{\bar{G}}_{m}$ is designed as:
$\mathbf{\bar{G}}_{m}=\textbf{y}^{pilot}_{m}\bm{\Psi},\ \ \in\mathbb{C}^{A\times L_{p}}$
where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as:
$\hat{\mathbf{g}}_{mt}=c_{mt}\mathbf{\bar{G}}_{m}\mathbf{e}_{i_{t}},$
where $\mathbf{e}_{i_{t}}$ is the $i_{t}$-th column of identity matrix $\mathbf{I}_{L_{p}}$.
II-B Uplink Transmission and Spectral Efficiency
Let $x_{t}$ denote the unit power uplink data signal transmitted by UE $t$. The uplink normalised data power corresponding to UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively. The uplink signal received at the $m$-th AP is:
$\textbf{y}^{u}_{m}=\sum\nolimits_{t=1}^{T}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
In decentralized data detection, unlike centralized approach where each AP sends received signal and channel estimates to CPU, each AP performs data detection for its served UEs using a local linear received combining vector. Following the proposed decentralized decoding architecture, each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR). These weights are applied to the locally detected data symbols before being sent to the CPU. This differs from LSFD, where local data and channel estimates are transmitted to the CPU for centralized processing. The locally weighted data symbol $\hat{x}_{mt}$ for UE $t$ at AP $m$ is given by
$\hat{x}_{mt}=a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m},$ (2)
where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for UE $t$ at AP $m$, and $a_{mt}$ is the corresponding local weight. These weighted estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
$\hat{x}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
After expanding $\textbf{y}^{u}_{m}$, the (3) can also be rewritten as
$\begin{split}\hat{x}_{t}&=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}\\ &+\sum_{\begin{subarray}{c}k=1\\ k\neq t\end{subarray}}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$ (4)
Here, the first term is the desired signal for UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term is the noise. Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the corresponding Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
Theorem 1.
[16, 13] : A lower bound on the uplink ergodic SE for UE $t$ is given by
$\textsc{SE}^{u}_{t}=L_{u}\log_{2}\big(1+\textsc{SINR}_{t}\big),$ (5)
where $L_{u}=\Big(\frac{1-\frac{L_{p}}{L_{c}}}{2}\Big)$ and $\textsc{SINR}_{t}$ is
$\textsc{SINR}_{t}=\frac{p^{u}_{t}|\textsc{DS}_{t}|^{2}}{\!\!\sum\limits_{k=1}^{T}\!\!p^{u}_{k}\mathbb{E}\left\{|\textsc{UI}_{tk}|^{2}\right\}-p^{u}_{t}|\textsc{DS}_{t}|^{2}+\mathbb{E}\left\{|\textsc{GN}_{t}|^{2}\right\}},$ (6)
where,
$\begin{split}&\textsc{DS}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\},\\ &\textsc{UI}_{tk}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk},\\ &\textsc{GN}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$
The SE lower bound, in (5), is valid regardless of the combining scheme used.
III Adaptive Combining and Local Weights Analysis
In this section, we first derive closed-form expressions for the uplink SE and local combining weights for both G-PFZF and G-PWPFZF schemes. We also formulate an optimization problem to partition the pilots into strong and weak.
III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
In [13], the combining vectors for UE $t$ is dependent on $\beta_{mt}$, which is unique to each UE. As a result, generating a distinct combining vector for every UE entails a number of multiplications that scales linearly with the total number of UEs. To reduce computational cost, we propose G-PFZF combining scheme, where each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot. The design of combining vectors depends on whether a pilot is strong or weak. Thus, each AP categorizes the pilots into two groups, strong and weak based on the local sum SE optimization. Unlike traditional PFZF, G-PFZF does not force the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
Remark 2: In our proposed combining schemes, a strong pilot indicates that a ZF combining vector is assigned to that pilot (and its corresponding UEs), while a weak pilot is assigned an MR combining vector.
For AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$. Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}{=}\{r_{m,1},r_{m,2},{\dots},r_{m,L_{\mathcal{S}_{m}}}\}$. The matrix $\mathbf{\bar{G}}_{m}{\in}\mathbb{C}^{A\times L_{p}}$ contains the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilot sequences, whereas the strong pilots utilize only $L_{\mathcal{S}_{m}}$ of them. To isolate the relevant columns of $\mathbf{\bar{G}}_{m}$, we define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}{=}[\mathbf{e}_{r_{m,1}},{\dots},\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]{\in}\mathbb{C}^{L_{p}\times L_{\mathcal{S}_{m}}}$, where $\mathbf{e}_{r}$ denotes the $r$-th column of the identity matrix $\mathbf{I}_{L_{p}}$. The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots. For a given pilot $i{\in}\mathcal{S}_{m}$, let $j_{mi}{\in}\{1,2,\dots,L_{\mathcal{S}_{m}}\}$ denote the index of pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}{\in}\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}{=}\mathbf{e}_{i}$.
Following the zero-forcing principle, the pseudo-inverse of this matrix can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$. To obtain the combining vector associated with a specific pilot, this pseudo-inverse term is multiplied by $\bm{\varepsilon}_{j_{mi}}$ and subsequently normalized. Thus, the local ZF combining vector for pilot $i{\in}\mathcal{S}_{m}$ at AP $m$ is given by:
$\mathbf{v}_{mi}^{\text{LZF}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}}{\sqrt{\mathbb{E}\big\{\big\|\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\big\|^{2}\big\}}},$ (7)
Similarly, the MR combining vector for pilot $i\in\mathcal{W}_{m}$ at AP $m$ corresponds directly to the channel vector associated with that pilot and is expressed as:
$\mathbf{v}_{mi}^{\text{MR}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{\mathbb{E}\big\{\big\|\mathbf{\bar{G}}_{m}\mathbf{e}_{i}\big\|^{2}\big\}}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{A\theta_{mi}}}.$ (8)
Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix ($\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$) with degree of freedom satisfying $A\geq L_{\mathcal{S}_{m}}+1$, can be rewritten as
${\mathbb{E}\Big\{\bm{\varepsilon}^{H}_{j_{mi}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\}}=\frac{1}{{(A-L_{\mathcal{S}_{m}})\theta_{mi}}},$ (9)
Also, the G-PFZF scheme suppresses the interference among the UEs whose pilot lies in $\mathcal{S}_{m}$ by sacrificing $L_{\mathcal{S}_{m}}$ degrees of freedom from total $A$ degrees to boost the desired signal. Therefore, for any UE $t$ such that $i_{t}\in\mathcal{S}_{m}$
$(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t},\end{cases}$ (10)
$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
Similarly, for any UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}{\mathbf{\hat{g}}_{mk}}\}=\begin{cases}\sqrt{A\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (12)
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (13)
Note 1: From (11) and (13), it is evident that local ZF achieves superior interference suppression compared to MR, since each AP can eliminate interference among UEs with pilots grouped as strong. However, this interference mitigation comes at the cost of $L_{\mathcal{S}_{m}}$ degrees of freedom, which reduces the array gain available for desired signal enhancement.
By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by
[…]
where $\mathcal{P}_{i_{t}}$ represents the subset of UEs sharing pilot $i_{t}$ and $\delta_{mi_{t}}$ is defined as:
$\delta_{mi_{t}}=\begin{cases}1,&\text{if }\ i_{t}\in\mathcal{S}_{m},\\ 0,&\text{otherwise}.\end{cases}$ (15)
The derivation of the SINR expression is given in Appendix -A.
Local Weight Design for G-PFZF Scheme: A fundamental limitation of the conventional LSFD architecture is its reliance on centralized optimization [13]. The optimal LSFD weights are calculated at the CPU using knowledge of global channel statistics, which creates a significant fronthaul signaling overhead and computational burden, thus undermining the scalability of the network. In this work, we break this dependency by introducing a novel decentralized decoding architecture where each AP $m$ independently computes a local interference-suppressing weight $a_{mt}$ for each UE $t$, based solely on its local channel information. Since these weights are derived from the LSFCs, they evolve much slowly than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
The primary objectives of this local weight design are twofold: (1) to maximize the contribution of the local estimate to the final SINR at the CPU, and (2) to actively mitigate the dominant interference components, specifically, pilot contamination and inter-user interference, at their source, prior to transmission. This approach eliminates the need for the APs to send the channel estimates to the CPU and the CPU to perform any complex optimization, reducing its role to that of a simple aggregator that sums the pre-weighted estimates from all APs.
The main idea behind designing decoding weights is that it should amplify the data signal received from the APs with strong local desired signal component and suppress those dominated by interference. In LSFD design, weights are computed through coordination among the APs, specifically the coherent interference terms. In contrast, the proposed decentralized framework eliminates inter-AP coordination, requiring each AP to determine its weight independently. Accordingly, an AP with a higher local SINR is given a larger weight than the one with a lower SINR in order to enhance the global SINR. For simplicity, the local weight is chosen to be equal to the local SINR.
The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A. Accordingly, the closed-form of the local SINR or local weight with respect to AP $m$ of UE $t$ is given by:
$a_{mt}=\text{SINR}^{\text{G-PFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (16)
where
$b_{kt}^{m}=\sqrt{(A-\delta_{{m}i_{t}}L_{\mathcal{S}_{m}})\gamma_{{m}k}},$
${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
The numerator represents the desired signal strength for the desired UE $t$ at AP $m$, capturing the gain achieved through the chosen combining strategy. The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference due to channel estimation error and noise.
By applying these weight to its local estimate, each AP independently produces a soft decision that is already pre-optimized for local interference suppression. This design ensures that the signals forwarded to the CPU require no further large-scale fading decoding, thereby completely eliminating the associated fronthaul overhead for channel statistics and the computational cost of centralized weight calculation. The result is a radically simplified and highly scalable uplink processing architecture that retains the performance benefits of distributed APs design.
Adaptive Pilot Grouping for G-PFZF Combining: This subsection details the local optimization framework that each AP employs to independently partition its set of pilots into strong and weak groups. This classification dictates the appropriate combining strategy, G-PFZF for strong pilots or MR for weak ones, and crucially, generalizes to pure FZF or pure MR should the optimization deem them to be optimal.
The basis of this optimization is the local SINR expression given in (16). As $L_{\mathcal{S}m}{=}\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit as given in (17).
[…]
This expression reveals a fundamental trade-off governed by the size of the strong pilot set, $L_{\mathcal{S}m}$. Allocating a pilot to the strong group ( $\delta_{mi}{=}1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs. However, this comes at the cost of a reduced array gain, $(A-L_{\mathcal{S}_{m}})$, for all UEs served by the AP, which diminishes the desired signal power. Conversely, classifying a pilot as weak preserves the full array gain but forgoes any interference suppression for its UEs. This inherent trade-off between maximizing signal strength and minimizing interference makes the grouping problem non-trivial and renders a fixed, network-wide threshold highly suboptimal.
To navigate the above mentioned trade-off, AP $m$ independently solves a local optimization problem, which is formulated as
$\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log_{2}\left(1+\frac{S_{mt}}{I_{mt}}\right),$ (18a)
subject to: $\ \delta_{mi}\in[0,1],\ \forall i,$ (18b)
$A-1\geq\sum_{i=1}^{L_{p}}\delta_{mi},$ (18c)
where $\bm{\delta}_{m}=[\delta_{m1},\delta_{m2},\cdots,\delta_{mL_{p}}]^{\intercal}$. The constraint (18c) represents the necessary condition of Wishart matrix for interference suppression. Also, intuitively, this would otherwise lead to invalid negative values for the desired signal power and coherent interference in the SINR expression. The optimization problem in (18) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost. To develop a more tractable formulation, we begin by relaxing the binary constraint (18b) into a continuous one:
$0\leq\delta_{mi}\leq 1,\ \forall i.$ (19)
To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint:
$\delta_{mi}-\delta^{2}_{mi}\leq 0,\ \forall i.$ (20)
The combined constraints (19) and (20) are equivalent to the original binary constraint (18b), as the only values between 0 and 1 that satisfy $\delta_{mi}(1-\delta_{mi})\leq 0$ are precisely 0 and 1. Thus, the problem can be written as
$\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log_{2}\left(1+\frac{S_{mt}}{I_{mt}}\right),$ (21a)
[…]
This problem can be solved using non-convex solvers or the successive convex approximation optimization method. Both of these are computationally expensive for large network deployment as non-convex solvers. Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [22, 23]. We first reformulate the problem to be unconstrained except for the box constraint (19). The reformulated problem is:
$\underset{\bm{\delta}_{m}}{\max~}$ $f(\bm{\delta}_{m}),$ (22a)
[…]
where
$\begin{split}f(\bm{\delta}_{m})&=\!\!\sum\limits_{t=1}^{T}\log_{2}(1+\frac{S_{mt}}{I_{mt}})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\Bigg),\end{split}$
where $\chi$ is a penalty parameter, and $\lambda_{1}$ and $\lambda_{2}$ are penalty weights.
Lemma 1.
: For a sufficiently large penalty parameter $\chi^{*}$, the solution to Problem (22) is equivalent to the solution of Problem (21).
Proof: As $\chi\rightarrow+\infty$, the penalty terms $\chi\lambda_{1}$ and $\chi\lambda_{2}$ enforce that any constraint violation is driven to zero for the solution to remain finite. Given that the original problem’s feasible set is bounded (as argued in [24, Proposition 1]), there exists a finite $\chi^{*}$ such that the solutions coincide.
To solve Problem (22) using the PGA method at each AP independently, we follow the steps outlined in Algorithm 1. Since the objective function (22a) is smooth and differentiable, the proximal step reduces to a projection. We initialize $\bm{\delta}_{m}$ and update it along with the gradient ascent direction of $f(\bm{{\delta}}^{(j)}_{m})$ with step size $\alpha$, followed by projection onto $[0,1]$ to satisfy the box constraint (19):
$\begin{split}\bm{{\delta}}^{(j+1)}_{m}&=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).\end{split}$ (23)
The components of gradient $\nabla f(\bm{\delta}_{m})$ can be written as
$\begin{split}\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=&\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}-\chi\Bigg(\!2\lambda_{1}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)\Big(1-2\delta_{mi}\Big)\\ &+2\lambda_{2}\max\!\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\!\Bigg),\end{split}$ (24)
where
$\frac{\partial S_{mt}}{\partial\delta_{mi}}=\begin{cases}-p^{u}_{t}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{mt},&\text{if }\ i_{t}=i,\\ -p^{u}_{t}\delta_{mi_{t}}\gamma_{mt},&\text{if }\ i_{t}\neq i,\end{cases}$ (25)
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\ \Big(1+\!\!\!\!\!\!\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\!\!\!\!\!\!\!\delta_{mj}\Big)\gamma_{m}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\gamma_{mk},\ \ \text{if }i_{t}=i,\\ \vskip 5.69054pt-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},\ \ \text{if }i_{t}\neq i,\end{cases}$ (26)
1:for all $m\in\mathcal{M}$ independently do
2:  Initialize: $j=1$, $k=1$, $\bm{\delta}_{m}^{(1)}$, $s^{(k)}=f(\bm{\delta}_{m}^{(1)})$, $\chi=1$, $\Delta>1$, $\epsilon=5e^{-3}$
3:  repeat
4:   repeat
5:    Compute:
$\bm{{\delta}}^{(j+1)}_{m}{=}\min\!\big(1,\max\!\big(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\big)\big)$
6:    Update $j=j+1$
7:   until $\Big|\frac{f(\bm{{\delta}}^{(j)}_{m})-f(\bm{{\delta}}^{(j-1)}_{m})}{f(\bm{{\delta}}^{(j-1)}_{m})}\leq\epsilon\Big|$
8:   Update $\chi=\chi\times\Delta$
9:   Update $k=k+1$
10:   Update $s^{(k)}=f(\bm{{\delta}}^{(j)}_{m})$
11:  until $\big|\frac{s^{(k)}-s^{(k-1)}}{s^{(k-1)}}\leq\epsilon\big|$
12:end for
Algorithm 1 Proximal Gradient Ascent (PGA) Method
Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proved in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure the convergence of the PGA method [22, 23]. In numerical simulations, small step size parameter $\alpha$ has been observed to achieve convergence for Algorithm 1.
At each AP, the computation cost of Algorithm 1 in each iteration depends upon the gradient computation step. The computation cost of calculating $f(\bm{\delta}_{m})$ is $\mathcal{O}(T^{2})$. Thus, the cost of calculating $\nabla f(\bm{{\delta}}_{m})$ is also $\mathcal{O}(T^{2})$. Since we have only $L_{p}$ variables, it converges in a few iterations only.
Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping as in [19, 13]. In UE-based grouping, each AP would have $T$ binary variables (one per UE), leading to $TM$ total variables across the network. In the proposed pilot-based grouping, each AP has $L_{p}$ binary variables, leading to $ML_{p}$ total variables. Since $L_{p}\ll T$ in dense deployments, this significantly reduces the computational cost.
III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference from other UEs that are also on strong pilots. This intra-group suppression leaves UEs on weak pilots vulnerable to dominant interference from the strong UEs.
The proposed G-PWPFZF scheme introduces a more comprehensive, protective strategy. In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that have a strong pilot. This is achieved by projecting the combining vectors for weak UEs onto the orthogonal complement of the subspace spanned by the strong UEs’ effective channels. Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, improving performance for users on weak pilots, at the cost of higher computation in comparison to G-PFZF scheme due to projection matrix computation at the AP.
In the G-PWPFZF scheme, the strong pilot $i{\in}\mathcal{S}_{m}$ are assigned with the combining vectors ${\mathbf{v}}^{\text{LFZ}}_{mi}$ as in (7). The combining vector allocated to pilot $i{\in}\mathcal{W}_{m}$ at AP $m$ is the projected MR (PMR) combining vector, and is given by:
${\mathbf{v}}^{\text{PMR}}_{mi}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi}}}\mathbf{B}_{m}\mathbf{\bar{G}}_{m}e_{i},$ (27)
where $\mathbf{B}_{m}$ is a matrix that projects the received signal onto the orthogonal complement of the subspace spanned by the channels of all strong pilots. It is defined as [19]:
$\mathbf{B}_{m}=\mathbf{I}_{A}-\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m},$
This projection ensures that the combiner for a weak pilot is orthogonal to the channel estimates of all strong-pilot UEs, thereby nullifying the dominant interference from that group.
The expected value of the effective channel gain for UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}{\mathbf{\hat{g}}_{mk}}\}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (28)
Furthermore
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\neq i_{t},\end{cases}$ (29)
These equations confirm that the projected MR (PMR) combiner successfully nullifies interference from all UEs sharing strong pilots $i_{k}\in\mathcal{S}_{m}$. This design effectively mitigates the most significant intra-network interference, enhancing the performance of the UEs with weaker channel conditions.
By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by
[…]
The derivation of the SINR expression is given in Appendix -C.
Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully decentralized manner. Here too the local weights are the same as the local SINR.
The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -C. Accordingly, the closed-form of the local SINR or local weight is given by:
$a_{mt}=\text{SINR}^{\text{G-PWPFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (31)
where
$b_{kt}^{m}=\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{{m}k}},$
${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
By employing these locally computed weights, each AP independently scales its data estimates to prioritize signals with high desired gain and low interference. This process effectively embeds interference suppression functionality at the source, thereby completely eliminating the need for centralized LSFD at the CPU and the associated fronthaul overhead for channel statistics.
Adaptive Pilot Grouping for G-PWPFZF Combining: This subsection discusses the local optimization framework for pilot grouping in the G-PWPFZF scheme. The objective and constraint structure remain identical to those of the G-PFZF scheme. The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping. The local SINR for UE $t$ at AP $m$ for G-PWPFZF scheme is given in (31). After substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ in (31):
$S_{mt}=p^{u}_{t}\Big(A-\sum\limits_{i=1}^{L_{p}}\delta_{mi}\Big)\gamma_{mt}$
$I_{mt}=\!\!\!\!\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!p^{u}_{k}\Big(\!A-\!\!\sum\limits_{i=1}^{L_{p}}\delta_{mi}\!\Big)\gamma_{mk}+\!\sum\limits_{k=1}^{K}\!p^{u}_{k}(\beta_{mk}{-}\delta_{mi_{k}}\gamma_{mk}){+}1.$
Consequently, the optimization algorithm (Algorithm 1) and the overall structure of the gradient $\nabla f(\bm{\delta}_{m})$ remain unchanged. The impact of the different SINR is confined solely to the calculation of the partial derivatives $\frac{\partial S_{mt}}{\partial\delta_{mi}}$ and $\frac{\partial I_{mt}}{\partial\delta_{mi}}$ within the gradient. For the G-PWPFZF scheme, these derivatives are:
$\frac{\partial S_{mt}}{\partial\delta_{mi}}=-p^{u}_{t}\gamma_{mt},$ (32)
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=-\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\gamma_{mk}-\sum\limits_{k\in\mathcal{P}_{i}}p^{u}_{k}\gamma_{mk}.$ (33)
These expressions replace their G-PFZF counterparts in the gradient calculation step of Algorithm 1. All other steps of the algorithm proceed identically.
IV Performance Evaluation
TABLE I: Baseline Simulation Parameters
Parameter Value
Number of APs ($M$) 100
Antennas per AP ($A$) 8
Number of UEs ($T$) 100
Channel bandwidth ($B$) 20 MHz
Uplink pilot symbols ($L_{p}$) 7
Coherence block length ($L_{c}$) 200 symbols
Maximum UE transmit power ($p_{\text{max}}$) 100 mW
Shadow fading standard deviation 8 dB
This section demonstrates the gains of the our proposed decentralized decoding and combining schemes in terms of the sum SE, per user SE and computation cost. We consider a network with $M$ APs and $T$ UEs, uniformly distributed in a square area of size $1\times 1$ km2. The relevant system parameters are mentioned in Table I. The LSFCs are generated according to the model in [3], incorporating shadow fading. Pilots are assigned using the scalable pilot assignment strategy from [3], ensuring efficient reuse under limited pilot resources. All APs are assumed to operate at full transmit power during both pilot training and data transmission. For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, are identified using numerical simulation, to provide the best performance. To ensure statistical reliability, all reported results are averaged over 500 independent network realizations.
IV-A Computation cost and Fronthaul cost
In this subsection, we analyze the computational and fronthaul signaling costs of our proposed decentralized decoding and combining schemes. The costs are computed following the approach outlined in [16, 19, 15, 10], where only multiplication and division operations are counted per coherence block, as these dominate the computational load.
TABLE II: Decoding Computation and Fronthaul Overhead per UE per Coherence Block
Decoding Type Fronthaul Cost Computation Cost
LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
Proposed Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
Table II compares the fronthaul and computational costs for optimal LSFD and the proposed decentralized decoding architecture. In conventional LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block. Furthermore, LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale network deployments.
The proposed decentralized decoding architecture eliminates these bottlenecks by avoiding centralized LSFD entirely. Each AP independently computes and applies interference-suppressing local weights, reducing the fronthaul load to only $L_{u}M$ complex scalars. The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions required in the LSFD architecture due to centralized AP coordination. This combination of reduced fronthaul overhead and scalable computation enables practical deployment in dense network scenarios.
Figure 2: Comparison of computational and fronthaul costs for different decoding and combining schemes.
As demonstrated in Fig. 2a, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to LSFD, particularly as the number of UEs increases. Similarly, Fig. 2b shows that eliminating channel estimate transmission reduces fronthaul overhead.
TABLE III: Computational cost per AP per Coherence Block
Scheme Combining Vector Computation
PFZF[19] $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AT$
G-PFZF $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AL_{p}$
PWPFZF[19] $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AT$
G-PWPFZF $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AL_{p}$
The computational cost per AP per coherence block for calculating the combining vectors is detailed in Table III. The primary distinction between the traditional and proposed generalized schemes lies in their scaling behavior relative to the number of users. The traditional PFZF and PWPFZF schemes exhibit computational costs that scale linearly with the total number of user equipments, $T$, as in the $AT$ term for PFZF and PWPFZF. This dependence creates a significant computational bottleneck in densed networks with massive connectivity. In contrast, the proposed G-PFZF and G-PWPFZF schemes achieve a fundamental scalability advantage by scaling with the number of pilots, $L_{p}$, rather than the number of UEs. This is reflected in the $AL_{p}$ term for G-PFZF and G-PWPFZF, which is a fixed system parameter independent of the user population. Fig. 2c further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.
IV-B Performance Evaluation
Figure 3: Distribution of strong pilot decisions.
Fig. 3 shows the distribution of strong pilot decisions ($L_{\mathcal{S}_{m}}$) across all APs. The baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}{=}6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions. In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF peaks at 2-4 strong pilots while G-PWPFZF shows an even more conservative distribution peaking at 1-3. This efficient allocation preserves spatial resources for signal enhancement. Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}{=}0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions. This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baselines.
Figure 4: The uplink sum SE comparison for various combining schemes.
Fig. 4 depicts the sum SE versus the number of UEs ($T$) for various combining schemes. The proposed decentralized optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the gaps widening as the number of UEs increases. Better performance of the proposed schemes highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP. In contrast, the proposed framework introduces a decentralized optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy. This allows an AP to adaptively select none, a subset, or all of the pilots as strong, based on its local channel conditions and specific geometric configuration. This per-AP flexibility enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement. Consequently, the proposed G-PWPFZF scheme achieves better performance over the G-PFZF scheme by suppressing interference of the weak UEs from the strong UEs at the APs.
Figure 5: The uplink sum SE comparison for various combining schemes.
Fig. 5 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the improved performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts. The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 6-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases. This expanding gap highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes. The baseline schemes, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}{>}A$), which leads to inefficient interference management. In contrast, our proposed decentralized framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring that the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}{<}A$). This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap with the number of pilots.
Figure 6: The uplink sum SE comparison for various combining schemes.
Fig. 6 illustrates the sum SE versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework. The G-PFZF and G-PWPFZF schemes demonstrate consistent superiority across all antenna configurations, with the performance gap exhibiting a distinctive trajectory that reflects adaptive resource utilization. At the minimal antenna configuration (2 antennas), where spatial degrees of freedom $A-L_{\mathcal{S}_{m}}$ are severely constrained, both proposed and baseline schemes face fundamental challenges in interference suppression. Nevertheless, our adaptive grouping strategy achieves measurable gains of 2.5% for G-PFZF and 6% for G-PWPFZF through an adaptive selection of combining schemes, dynamically switching between the G-PFZF/G-PWPFZF, and the MR at each AP based on local conditions. In the mid-range antenna configuration (4-8 antennas), where careful trade-off between interference suppression and signal enhancement becomes crucial, our proposed method demonstrates performance gain of 10-14% for G-PFZF and 10-18% for G-PWPFZF, showcasing their superior grouping capabilities. As antenna numbers increase further (10-16 antennas), providing abundant spatial resources that benefit all schemes, the performance gap narrows to 4.5-6.5% for G-PFZF and 3.5-5.5% for G-PWPFZF, though our approach maintains consistent superiority.
Figure 7: The uplink sum SE comparison for various combining schemes.
Fig. 7 illustrates the sum SE versus the number of UEs ($T$), comparing local and LSFD-based combining schemes. For our G-PFZF and G-PWPFZF schemes, the performance gap between local weights and optimal LSFD remains minimal at just 3% across all UE densities. This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and reducing fronthaul overhead by eliminating the need to share channel state information across the network. Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with LSFD by margins of 4.5-9% and 3.5-7.5% respectively. This demonstrates that the gains from our adaptive per-AP grouping strategy are so substantial that they outperform the benefits of optimal coordination applied to suboptimal threshold-based grouping. These results confirm that our fully decentralized approach achieves a better balance between performance and practicality.
Figure 8: The uplink sum SE comparison for various combining schemes.
Fig. 8 illustrates the sum SE versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes. The performance gap between local and LSFD-based weights is less than 5.5% for sparse pilot configurations, where high pilot contamination makes centralized LSFD better equipped to handle interference. However, as the number of pilots increases to moderate and high values, the gap reduces to less than 1%, demonstrating near-identical performance. This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination. Notably, even in the worst-case scenario with limited pilots, the performance penalty remains modest, validating that the substantial fronthaul and complexity reductions of our decentralized decoding approach come with acceptable performance trade-offs across all pilot sequences. Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
Figure 9: The uplink SE per user comparison for various combining schemes.
Fig. 9 presents the 90%-likely per-user SE across different combining schemes, demonstrating significant improvements in user fairness through our decentralized optimization framework. The proposed G-PFZF scheme achieves 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF. These gains confirm that our adaptive per-AP grouping strategy not only enhances the overall system capacity but fundamentally transforms user experience across the network. The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users.
Figure 10: The uplink SE per user comparison for various decoding schemes.
Fig. 10 presents the 90%-likely per-user SE comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our decentralized framework. While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving fully decentralized architecture, substantially reduced fronthaul overhead, and lower computational cost. Noticeably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%. These results demonstrate that the performance gains from our adaptive per-AP grouping strategy are so significant that they outweigh the benefits of optimal LSFD coordination applied to suboptimal threshold-based grouping.
V Conclusion and Future Work
This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in D-mMIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs. Through the proposed local ZF framework, we have demonstrated that enabling each AP to independently determine its combining strategy via local optimization, classifying pilots as strong or weak without fixed thresholds and dynamically switching among PFZF/PWPFZF, FZF, or MR, yields significant performance improvements. The resulting generalized schemes, G-PFZF and G-PWPFZF, consistently outperform their fixed-threshold counterparts across all evaluated scenarios, achieving substantial gains in both sum and per user spectral efficiency. Moreover, the introduction of pilot-dependent combining vectors and interference-suppressing local weights applied distributively at each AP eliminates the need for centralized LSFD, thereby drastically reducing fronthaul overhead and computational cost. Remarkably, this decentralized approach incurs only a small performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD. These findings collectively establish that proposed decentralized decoding in the network, not centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems without compromising on quality-of-service.
In future, we plan to investigate the robustness of the proposed decentralized architecture and generalized combining schemes under more practical and challenging conditions, such as hardware impairments.
-A Derivation of closed-form SINR expression for G-PFZF combining
In order to derive the closed-form expression, we need few expectations computation or expansions
$\begin{split}\mathbb{E}\Big\{\Big|\!\!\sum_{m}\!\!\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}\!\!=&\!\sum_{m}\!\mathbb{E}\Big\{\!\Big|\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}+\!\Big|\!\sum_{m}\!\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\Big\}\Big|^{2},\end{split}$ (A1)
$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Big\}=$ $\beta_{mk}-\gamma_{mk}$ (A2)
$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}{\mathbf{n}}_{m}\Big|^{2}\Big\}=$ $1.$ (A3)
The desired signal term in (6) can be written as
$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2},$
where $\mathcal{Z}_{i_{t}}$ and $\mathcal{Y}_{i_{t}}$ denote the subsets of APs where pilot $i_{t}$ is classified as strong and weak, respectively. Using (10) and (12), the $|\text{DS}_{t}|^{2}$ can be written as:
$\begin{split}|\text{DS}_{t}|^{2}&=\!\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\\ &=\Big|\sum_{m=1}^{M}\!a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2},\end{split}$ (A4)
The interference term $\mathbb{E}\big\{|\text{UI}_{tk}|^{2}\big\}$ in (6) can be expanded as:
$\begin{split}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}&=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\!+\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\\ &+2\mathbb{E}\Bigg\{\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}.\end{split}$ (A5)
Using (A1), the first term ($A_{1}$) of (A5) can be rewritten as:
$\begin{split}A_{1}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mk}+\tilde{\mathbf{g}}_{mk})\Big|^{2}\!\Bigg\}=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\!+\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\\ &=\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using (10), (11) and (A2), we can rewrite $A_{1}$ as
$\begin{split}A_{1}&=\Big|\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A6)
Using (A1), the second term ($A_{2}$) of (A5) is expanded as:
$\begin{split}A_{2}&=\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using the (12), (13) and (A2), we can rewrite $A_{2}$ as
$\begin{split}A_{2}&=\Big|\!\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\!\!+\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}\beta_{mk}.\end{split}$ (A7)
Finally, the third term ($A_{3}$) of (A5) can be written as:
$\begin{split}A_{3}=2\Big(\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mt}}\!\Big)\Big(\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big).\end{split}$ (A8)
Using (A6), (A7) and (A8)
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\!\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A9)
Using (A3), the term $\mathbb{E}\big\{|\text{GN}_{t}|^{2}\big\}$ of interference in (6) can be written as:
$\begin{split}&\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}+\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (A10)
Using the (A4), (A9) and (A10) in (6) gives the SINR closed-form expression in (14).
-B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$
To establish the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, we analyze its component functions and their boundedness over the compact domain $\bm{\delta}_{m}\in[0,1]^{L_{p}}$.
The gradient component in (24) is $g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}$ bounded as
  • $0<S_{mt},I_{mt}\leq C_{1}$ (always bounded)
  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded from (25) and (26))
  • $\max(0,\delta_{mi}-\delta^{2}_{mi})\in[0,\frac{1}{4}]$ (bounded on $[0,1]$)
  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$
  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1),\qquad C_{3}\leq\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{4}$,
where $C_{1}$, $C_{3}$, $C_{3}$ and $C_{4}$ are finite constants.
The partial derivatives of $g(\delta_{mi})$ is given by
$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\!\!\sum_{t=1}^{T}\!\!\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}+\frac{I_{mt}}{I_{mt}+S_{mt}}\Bigg(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}\\ &-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\Bigg)\Bigg]-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}\\ &-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
For all $\delta_{mi}\in[0,1]$, since all component functions $\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}$ are bounded and continuously differentiable on the compact domain $0,1]^{L_{p}}$, the Hessian matrix $\nabla^{2}f(\bm{\delta}_{m})=\nabla g(\bm{\delta}_{m})$ exists and is continuous. Applying the mean value theorem for vector-valued functions, for any $\mathbf{x},\mathbf{y}\in[0,1]^{L_{p}}$, there exists $\mathbf{z}$ on the line segment between $\mathbf{x}$ and $\mathbf{y}$ such that:
$\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\|\leq\|\nabla^{2}f(\mathbf{z})\|_{2}\|\mathbf{x}-\mathbf{y}\|\leq L\|\mathbf{x}-\mathbf{y}\|$
This establishes the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, ensuring convergence of the proximal gradient algorithm.
-C Derivation of closed-form SINR expression for G-PWPFZF combining
The derivation of G-PWPFZF SINR closed-form expression follows the same step as the derivation of G-PFZF SINR closed-form expression, the only difference is that instead of ${\mathbf{v}}^{\text{MR}}_{mi_{t}}$, we have ${\mathbf{v}}^{\text{PMR}}_{mi_{t}}$ combining vector.
The desired signal term in (6)
$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2}.$
Using (10) and (28), then
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (C1)
Following the steps in Appendix -A, the first term in (6) can be written as
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (C2)
Similarly, the last term of interference in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (C3)
Using the (C1), (C2) and (C3) in (6) gives the SINR closed-form expression in (30).